OpenStudy (anonymous):

5/6f - 3/4f + 3/4 = 1/2 find f

OpenStudy (anonymous):

subtract the f terms and combine the constant terms 5/6 - 3/4 = 1/12 1/2 - 3/4 = -1/4 1/12f = -1/4 multiply by 12 to make it easier f=-3